# Causality, Emergence, Causal Emergence, Causal Emergence Identification, Causal Representation Learning, World Model

### 28 Dec 2023 v1

@@ 3.5.3 About Effective Information
Effective information offers a solution to the aforementioned limitation of mutual information. Given that TPM can capture the inherent nature of a system, effective information is likewise an intrinsic property of the system.
Causal emergence involves assuming a uniform distribution for the input data, which can be seen as a way of manipulating the input data to isolate the interference of the original distribution of input data. This manipulation is formalized using Judea Pearl’s causation theory, specifically through the "do" operation. By imposing this assumption of a uniform distribution, we create a controlled environment with equal probabilities of $X_{t}$’s states and allows us to better understand the causal relationships between $X_{t}$ and $X_{t+1}$. Therefore, the term "causal emergence" is used.
We can calculate EIs for both TPMs of macro- and micro- dynamics, and their difference is defined as the measure of causal emergence, that is:
$CE=EI(TPM_{M})-EI(TPM_{m}),$ (9)
where $TPM_{M}$ represents the TPM of macro-level dynamics, and $TPM_{m}$ represents the one of micro-level. $CE$ measures the degree of causal emergence. If $CE>0$, then causal emergence occurs, otherwise, does not occur.
However, there is a limitation for EI and CE when comparing two dynamics that differ significantly in size. This is because the value of effective information relies on the number of possible states within the system, denoted as $N$, with an upper bound of $\log_{2}(N)$. To facilitate the comparisons between different coarse-graining strategies and scales, effective information is often normalized resulting in a metric known as the effect coefficient $Eff$.
$Eff=\frac{EI}{\log_{2}(N)}$ (10)
The value of effect coefficient is always between 00 and $1$, representing the proportion of effective information being transferred from current states to future states. If the information is fully transferred, the effect coefficient is $1$.
In addition to characterizing causal emergence, effect coefficient can be further broken down into two meaningful components: "Determinism," which represents the certainty of a current state evolving into a certain state, or diverging into multiple states in the next moment, and "Degeneracy," which represents the possibility of multiple current states converging into one state in the next moment.
$Eff=Determinism-Degeneracy.$ (11)
Figure 6: Examples of effect coefficient, determinism, degeneration, and how coarse-graning can change them. (a) Deterministic and zero degenerative system; (b) Deterministic system with zero effect coefficient due to high degeneracy; (c) Partially deterministic system; (d) Deterministic system with low effect coefficient; (e) Coarse-grained system from c or d, resulting in a fully deterministic system with high effect coefficient.
In Figure 6, the square cells represent for the elements of TPM matrix, and the grey scale represents for the value of TPM elements. Example (a) is a bijective system, meaning that all information from current states can be transferred to future states without loss. It is fully deterministic with zero degeneracy. Example (b) is an extreme case where all current states lead to only one future state, illustrating that high determinism does not necessarily imply a high effect coefficient. Systems (c) and (d) differ but have the same effect coefficients. Finally, system (e) is a coarse-grained version of either system (c) or (d), demonstrating two important points: first, different microscopic systems can be coarse-grained to the same macroscopic system, and second, causal emergence can be quantitatively captured from coarse-graining by the increased effect coefficient.
Despite the fact that effective information can identify and quantify causal emergence, as previously mentioned, there are other independently developed measures of causation, such as Humean constant conjunction Hume (2016), Eells’s measure of causation as probability raising Eells (1991), Suppes’s measure of causation as probability raising Suppes (1973), and Judea Pearl’s measures of causation Pearl (2009) etc. Hoel found all these measures, including effective information can be related to a small set of "causal primatives" Comolatti and Hoel (2022): the probability representations of sufficiency and necessity (and its variant) of cause $c$ and effect $e$.
$suff(e,c)=P(e|c)$ (12)
$nec^{\dagger}(e,c)=P(e|C)$ (13)
$-\𝑐nec(e,c)=1-P(e|C\verb|\|c)$ (14)
Where $C$ stands for the set of all causes and $-\𝑐C\verb|\|c$ stands for the set of all causes but $c$. Hoel pointed out that determinism is actually a measure of sufficiency, while degeneracy is an inverse measure of necessity. The consistency in causal primitives indicates there is a scientific foundation of these individually developed measures of causation.
One limitation of EI is its global nature, as highlighted by the fact that the summation of additional terms encompasses the entire state space, as defined in Equation 8. Therefore, Varley (2022) proposed the conception of flickering emergence to decompose EI by each term being added in Equation 8. This new feature can characterize the local properties of a Markov dynamics.
3.5.4 Examples of Causal Emergence
The example above demonstrates a coarse-graining strategy that is intuitive, as it involves aggregation at the state level. However, in reality, coarse-graining often occurs among variables, and the combination of these variables produces different states. Let’s look at a more complicated example of a boolean network that given in Hoel et al. (2013) and we will inspect the causal emergence when coarse-graining its state transition mechanism. A, B, C and D are 4 binary variables whose state transational relations are as shown in Figure  7(a).
Figure 7: The mechanism and transition probability matrix of a boolean network. Each variable’s next state is determined by the combination of the current states of two other variables, as shown in the boolean network (a). If we examine an atomic subsystem with two inputs and one output, such as AB -> C, the next state of C depends on the present states of A and B. If A=0 and B=0, then the probability of C being 1 is 70%, while the probability of it being 0 is 30%, as indicated in table (b). This same microscopic dynamics applies to other atomic subsystems such as AB -> D, CD -> A, and CD -> B. When we expand to a subsystem of two inputs and two outputs, AB -> CD and CD -> AB, we obtain the transition probability matrix (TPM) demonstrated in table (c). Finally, we extend to the dynamics of the entire system ABCD -> ABCD, whose TPM is represented in a heatmap (d). The effective information and relative metrics of the entire system are as below: $EI_{mi}=1.1486$, $Eff=0.2871$, $Determinism=0.3390$, $Degeneracy=0.0519$
[…]
Figure 8: An example to show how different coarse-graining strategies on the same boolean network shown in Figure 7 can lead to different boolean dynamics. (a) Coarse-graining strategy 1. (b) Macro-dynamics leading by strategy 1. (c) The corresponding heatmap of TPM and effective information metrics of the macroscopic system by method 1. (d) Coarse-graining strategy 2. (e) and (f) are the macro-dynamics and TPM correspond to the coarse-graining strategy 2 in (d) respectively.
The effective information values for the two macroscopic systems are $EI_{ma1}=1.55$ and $EI_{ma2}=0.05$, respectively. It is evident that the effective information has increased in the first coarse-graining scenario ($EI_{ma1}(1.55)>EI_{mi}(1.15)$), while it has decreased in the second case ($EI_{ma2}(0.05)<EI_{mi}(1.15)$). Thus, causal emergence is observed in the first coarse-graining instance. This implies that 1) proper coarse-graining can lead to causal emergence in a macroscopic system, and 2) not all coarse-graining approaches can result in higher effective information. This example illustrates that coarse-graining can play an important role in emergence, and effective information is a well defined quantification method. This gives us the opportunity to identify causal emergence in a given system. There is a question coming next: what coarse-graining strategy can generate the maximum effective information? We will address this question in section 4.
3.6 Extensions and Applications
We will discuss the extensions and applications of Hoel’s theory of causal emergence in this sub-section.
3.6.1 Extensions on Continuous Systems
Effective information and its associated metrics have been shown to be useful in quantifying causal emergence. However, there are some limitations to be aware of. First, the method is only applicable to discrete-state systems. Second, knowledge of the system’s transition probability matrix (TPM) is required to calculate effective information metrics. Efforts have been made to apply or extend the concepts of effective information to continuous systems, such as Chvykov and Hoel (2020); Varley (2020); Zhang and Liu (2022).
Ordinal Partition Network
One approach to incorporating effective information in continuous systems is to transform the continuous variables into discrete ones by ordinal partition network (OPN) method Varley (2020). The basic idea is to transform the time series data generated by a continuous dynamical system into sub-time series of dimensionality $d$ with discrete values. And then to establish a Transition Probability Matrix (TPM) between these discrete series. This approach involves three steps: firstly, sampling the original time series with a specified time interval $\tau$ to create sub time series of dimensionality $d$; secondly, ranking these sub-series based on their values; and finally, utilizing the $d$-dimensional vectors of rank orders to represent the original sub-series. By employing rank orders, this method ensures that the values are restricted to integers within a certain range.
In Varley (2020), the author applied OPN method on Rossler chaotic attractor system. By quantifying effective information, he noticed that effective information (such as determinism and degeneracy) is sensitive to the critical phase transition between period and chaos Varley (2020).
Causal Geometry
Alternatively, Pavel Chvykov and Erik Hoel extended the causal emergence framework to continuous systems and proposed the conception of causal geometry Chvykov and Hoel (2020).
To understand the basic idea of this work, we suppose the continuous system considered is a continuous functional map and assume that the uncertainties are small disturbances added to this deterministic function, that is:
$y=f(x)+\varepsilon,$ (15)
where $x$ is the input variable defined on the interval $[-L/2,L/2]$ and $L$ is a very large constant, $y$ is the output variable, and $\varepsilon$ is a gaussion noise with zero mean and standard deviation $\epsilon\ll L$. Then, an approximate form of the effective information for such continuous system can be obtained by Gaussian integral.
The reason why we set the domain of $x$ on $[-L/2,L/2]$ is to guarantee that the Gaussian integral can be implemented, and also facilitate to define a uniform distribution on it.
This form can be easily generalized to a dynamical system with discrete time steps once we replace $x$ and $y$ as $x(t)$ and $x(t+1)$ respectively. The formula also can be extended to higher dimension. Suppose $\mathbf{x}\in[-L/2,L/2]^{n}\subset\mathcal{R}^{n}$ and $\mathbf{y}\in\mathcal{R}^{m}$, where $n$ and $m$ are positive integers, Equation 16 can be generalized to the following form:
where, $\Sigma$ is the covariance matrix of the Gaussian noise $\varepsilon$, $U([-L,L]^{n})$ represents the uniform distribution on the hypercube $[-L,L]^{n}$, and ||$\cdot$|| ⋅ | is the operation to take absolute value, and $\det$ is to take the determinant.
Notice that the conditional distribution of $\mathbf{y}$ given $\mathbf{x}$ is a Gaussian distribution, that is $p(\mathbf{y}|\mathbf{x})=\mathcal{N}(f(\mathbf{x}),\Sigma)$, thus the term in the expectation in Equation 17 can be written as:
and this is the determinant of the negative Fisher information metric for the distribution $p(\mathbf{y}|\mathbf{x})$:
it measures the sensitivity of the logarithmic $p(\mathbf{y}|\mathbf{x})$ on $\mathbf{x}$. Where $\partial_{\mu}\equiv\frac{\partial}{\partial\mathbf{x}_{\mu}}$ represents the partial derivative with respect to the $\mu$th component of $\mathbf{x}$. Therefore, we have defined a distance metric for the Riemann manifold =$\mathcal{M}=\{p(\mathbf{y}|\mathbf{x})\}$ on the parameter space $\mathbf{x}\in[-L/2,L/2]^{n}$, and it is the collections of all possible distributions of $p(\mathbf{y}|\mathbf{x})$. This is where the term of “geometry” in this framework comes from.
Finally, we can obtain an expression for EI with Fisher information metric:
This formula may be generalized to the cases where $p(\mathbf{y}|\mathbf{x})$ is non-Gaussian distribution. Once the distribution function is known, we can obtain its Fisher information metric, then EI can be calculated via Equation 20. The reason behind is that the whole manifold $p(\mathbf{y}|\mathbf{x})$ for any $\mathbf{x}$ can be understood as a concatenation of local Gaussian distributions.
While, in Chvykov and Hoel (2020), the authors consider a more general case, in which the intervention noise is added on the input(intervention) variable $x$ in Equation 16. The intervention noise is $\xi\sim\mathcal{N}(0,\delta^{2})$, and $\delta$ is the standard deviation of $\xi$. While, $\varepsilon$ is called observational noise to differentiate with $\xi$, and it is added on the output(observational) variable $y$ according to Equation 16. Finally, if two kinds of noises are both considered, and if $L=1$ and $\epsilon\ll 1$, Equation 16 becomes:
[…]
This is the formula of EI for continuous mapping function (Equation 15) given by Chvykov and Hoel (2020). Using this equation and a logistic function for $f(x)$, the authors compared the EIs of a more continuous map (Equation 15) and a more discrete map with an adjustable parameter in the logistic function. They discovered that when the noise level is low, the continuous map may exhibit higher EI than the discrete map. However, as the noise level increases, discretizing the mapping function can lead to a model with higher EI. This phenomenon helps to explain why digital circuits eventually outperform analog circuits in mitigating noise interference: the binarization and coarse-graining strategy of digital circuits suppresses the propagation and diffusion of noise.
To generalize the information geometry to the case with intervention noise and observational noise, let us introduce a new intermediate variable $\theta\in\Theta\subset\mathcal{R}^{l}$ with dimensionality $l$ such that we can not control $\mathbf{y}$ by intervening $\mathbf{x}$ directly. Instead, we can intervene $\mathbf{x}$ to influence $\theta$, and to affect $\mathbf{y}$ indirectly. Therefore, the three variables form a Markov chain: $\mathbf{x}\rightarrow\theta\rightarrow\mathbf{y}$.
[…]
Finally, the EI calculation formula for the causal geometry is:
@@ 3.6.1 Extensions on Continuous Systems
in which, we have set $L=1$ and $m=l=n$ to reduce the number of free parameters, and $I_{n}$ is the identity matrix with size $n$, $V_{I}=\int_{\Theta}\sqrt{|\det(h_{\mu\nu})|}d^{l}\theta$.
@@
One of the application areas for continuous EI is on neural networks. In Marrow et al. (2020), the authors applied EI to analyze the causal effects for different layers of neural networks and compare the results with information bottleneck theory. The method to calculate EI that they used is to convert continuous variables into discrete ones by dividing the domains of input and output variables into small regions. However, this method has high computational complexity, and is hard to extend to high dimensions.
In Zhang and Liu (2022), the authors propose a new numeric method for calculating EI on neural networks. The basic idea behind this method is to understand a well-trained feed-forward neural network as a stochastic mapping following:
@@
where $\mathbf{x}\in\mathcal{R}^{n}$ is the input variable, and $\mathbf{y}\in\mathcal{R}^{m}$ is the output variable, $f:\mathcal{R}^{n}\rightarrow\mathcal{R}^{m}$ is the deterministic map of the neural network, $\varepsilon$ is a Gaussian noise with zero means and the covariance matrix $\Sigma=diag(\sigma_{1}^{2},\sigma_{2}^{2},\cdot\cdot\cdot,\sigma_{m}^{2})$, where $\sigma_{i}$ is the mean square error on dimension $i$ on the whole training data set. In this way, the neural network can be regarded as a Gaussian distribution $p(\mathbf{y}|\mathbf{x})=\mathcal{N}(f(\mathbf{x}),\Sigma)$.
Therefore, Equation 17 can be used on this neural network to calculate EI. In Zhang and Liu (2022), the authors use Monte Carlo integration method to estimate the expectation in Equation 17, this technique reduces the computational complexity in a large degree.
However, when the authors compared the results calculated by Equation 17 for different neural networks with different dimensions, they obtained unreasonable results: EI increases with the number of dimensions, thus, causal emergence is always not be observed because micro-dynamics has unreasonably larger EI. Another defect of Equation 17 is that the parameter $L$ always dominates the value of the expression which should be removed from the equation. One way to solve this issue is to use Eff instead of EI (see Equation 10), however, not all $L$s in the Equation 17 can be eliminated. Therefore, the authors invented a notion of dimension averaged EI,
$dEI\equiv=\frac{EI}{m}$ (24)
which simply divide EI by the dimension of the output variable $\mathbf{y}$. By using dimension averaged EI, $L$ can be also removed if we calculate dimensional averaged causal emergence for the neural networks $f_{M}$ and $f_{m}$ for macro- and micro- dynamics respectively, that is:
@@
where, dCE is the dimension averaged causal emergence measure. If we bring Equation 17 into Equation 25, we get:
@@
in which, $L$ is eliminated.
3.6.2 Causal emergence in complex network
Many complex systems can be represented by networks, therefore, to apply the framework of causal emergence on networks is necessary. However, two problems should be solved in advance to apply Hoel’s framework, one is to assign a dynamics on the studied network because networks are static structures without any dynamical properties, another one is to coarse-grain the network.
Klein and Hoel Klein and Hoel (2020) solved the first problem by considering random walks on complex networks. Subsequently, the TPM of the system is defined based on the transfer probabilities between nodes of a large number of random walkers. Due to the good mathematical properties of the random walk dynamics on graphs, the authors can establish an explicit expression of EI on networks, and the final form only relies on the weighted normalized adjacency matrix $w_{ij}$($\sum_{j}w_{ij}=1$) of the network. The expression is as follow:
@@ 3.6.2 Causal emergence in complex network
Where, $H(\langle W^{out}_{i}\rangle)$ represents the Shannon entropy calculated from the distribution of the averaged out-weights across all nodes ($W^{out}_{i}\equiv\{w_{ij}|j=1,2,...,N\}$), which characterizes the determinism of the random walk dynamics; $\langle H(W^{out}_{i})\rangle$ is the average Shannon entropy for all nodes which characterizes the degeneracy of the dynamics. And $H(W^{out}_{i})$ is the uncertainty of node $i$.
@@
To tackle the second problem, the greedy algorithm is used to group nodes to form a macro network in Klein and Hoel (2020). It is worth noting that when merging micro nodes into macro nodes, the TPM of the corresponding macro network can be derived by merging the probabilities in TPM of the micro-level. To ensure that the grouped macro network maintains the same random walk dynamics as the original one, dynamic consistency test is implemented.
@@
In the experimental section, the paper explores the application of greedy search for identifying causal emergence in complex networks. The experiments are conducted on artificial networks, as well as four types of real networks. In the case of ER random networks, the size of effective information is solely dependent on the connection probability $p$, and converges to $-\log_{2}p$ with the increase of network size. Additionally, a significant finding reveals the presence of a phase transition point, where the average degree of the network reaches approximately $\log_{2}N$. Beyond this point, the random network structure does not contain additional information with increasing size. In the case of preferential attachment (PA) networks, when $\alpha<1.0$ ($\alpha$ represents the degree of preferential attachment), the effective information of the network increases as the network size expands. Conversely, when $\alpha>1.0$, the opposite conclusion is reached. The scale-free network corresponding to $\alpha=1.0$ represents the critical boundary of growth. Regarding real networks, the authors find that biological networks exhibit the lowest EI due to the presence of significant noise, which can be removed by effective coarse-grained, and causal emergence is the most significant compared with other types of networks. However, technical networks exhibit sparsity and non-degeneracy, resulting in higher average efficiency and more specific node relationships. Consequently, they possess the highest EI among the networks studied.
@@
The network coarse-graining method mentioned above is the greedy algorithm. However, when the network is very large, the efficiency of this method remains considerably low. Following that, Griebenow et al. Griebenow et al. (2019) introduced a spectral clustering-based approach to identify causal emergence within networks. More specifically, the method involved performing eigenvalue decomposition of the TPM, followed by constructing a similarity matrix using the eigenvectors of the nodes. The OPTICS algorithm was employed to cluster the nodes, and nodes belonging to the same cluster were aggregated into a macro node. Subsequently, the maximum value of EI was selected by utilizing the linear search distance hyper-parameter $\epsilon$. In this paper, the author additionally proposed a gradient descent algorithm based on deep learning. This approach encompasses several steps, beginning with the random initialization of a grouping matrix. Then this matrix is used to construct a macro network by combining the micro network. Finally, the grouping strategies are automatically learned through the process of maximizing effective information within the macro network. However, this method often falls into the local optimal solution.

### 25 Feb 2024 v3

@@ 3.5.5 Extensions to Continuous Systems
Finally, the EI calculation formula for causal geometry is:
@@
where we have set $L=1$ and $m=l=n$ to reduce the number of free parameters, and $I_{n}$ is the identity matrix with size $n$, $V_{I}=\int_{\Theta}\sqrt{|\det(h_{\mu\nu})|}d^{l}\theta$.
@@
One of the application areas for continuous EI is neural networks. In Marrow et al. (2020), the authors applied EI to analyze the causal effects for different layers of neural networks and compared the results with the information bottleneck theory. The method they used to calculate EI was to convert continuous variables into discrete ones by dividing the domains of the input and output variables into small regions. However, this method has high computational complexity and is challenging to extend to high dimensions.
In Zhang and Liu (2022), the authors proposed a new numeric method for calculating EI on neural networks. The basic idea behind this method is to understand a well-trained feed-forward neural network as a stochastic mapping, following:
@@
where $\mathbf{x}\in\mathcal{R}^{n}$ is the input variable, $\mathbf{y}\in\mathcal{R}^{m}$ is the output variable, and $f:\mathcal{R}^{n}\rightarrow\mathcal{R}^{m}$ is the deterministic map of the neural network. Additionally, $\varepsilon$ is a Gaussian noise with a mean of zero and a covariance matrix $\Sigma=diag(\sigma_{1}^{2},\sigma_{2}^{2},\cdot\cdot\cdot,\sigma_{m}^{2})$, where $\sigma_{i}$ is the mean square error on dimension $i$ across the entire training dataset. In this way, the neural network can be regarded as a Gaussian distribution $p(\mathbf{y}|\mathbf{x})=\mathcal{N}(f(\mathbf{x}),\Sigma)$.
Therefore, Equation (20) can be used on this neural network to calculate EI. In Zhang and Liu (2022), the authors used the Monte Carlo integration method to estimate the expectation in Equation (20). This technique significantly reduces the computational complexity.
However, when the authors compared the results calculated using Equation (20) for various neural networks with different dimensions, they obtained unreasonable results: EI increased with the number of dimensions. Thus, causal emergence was not always observed because micro-dynamics exhibited unreasonably larger EI. Another drawback of Equation (20) is that the parameter $L$ always dominates the value of the expression and should be removed from the equation. One way to solve this issue is to use Eff instead of EI (see Equation (10)). However, not all $L$s in Equation (20) can be eliminated. Therefore, the authors introduced the notion of dimension-averaged EI, defined as
$dEI\equiv\frac{EI}{m}$ (27)
which simply divides EI by the dimension of the output variable $\mathbf{y}$. By using dimension-averaged EI, $L$ can be removed when calculating dimension-averaged causal emergence for the neural networks $f_{M}$ and $f_{m}$ representing macro- and micro-dynamics, respectively, given by
@@
where dCE is the dimension-averaged causal emergence measure. If we incorporate Equation (20) into Equation (28), we obtain:
@@
in which $L$ is eliminated.
3.5.6 Causal Emergence in Complex Networks
Many complex systems can be represented by networks; therefore, applying the framework for causal emergence to networks is necessary. However, two problems should be solved in advance to apply Hoel’s framework: one is to assign dynamics to the studied network because networks are static structures without any dynamical properties and the other is to coarse-grain the network.
Klein and Hoel Klein and Hoel (2020) addressed the first problem by considering random walks on complex networks. Subsequently, the TPM of the system was defined based on the transfer probabilities between nodes of a large number of random walkers. Due to the good mathematical properties of the random walk dynamics on graphs, the authors established an explicit expression for EI on networks, and the final form only relied on the weighted normalized adjacency matrix $w_{ij}$($\sum_{j}w_{ij}=1$) of the network. The expression is as follows:
@@ 3.5.6 Causal Emergence in Complex Networks
where $H(\langle W^{out}_{i}\rangle)$ represents the Shannon entropy calculated from the distribution of the averaged out-weights across all nodes ($W^{out}_{i}\equiv\{w_{ij}|j=1,2,\dots,N\}$), which characterizes the determinism of the random walk dynamics; $\langle H(W^{out}_{i})\rangle$ is the average Shannon entropy for all nodes, which characterizes the degeneracy of the dynamics; and $H(W^{out}_{i})$ is the uncertainty of node $i$.
@@
To address the second problem, in Klein and Hoel (2020), the greedy algorithm was used to group nodes to form a macro-network. It is worth noting that when merging micro-nodes into macro-nodes, the TPM of the resulting macro-network can be derived by merging the probabilities from the TPM of the micro-level. To ensure that the grouped macro-network maintains the same random walk dynamics as the original one, a dynamic consistency test is implemented.
@@
In the experimental section, the aforementioned paper explored the application of a greedy search for identifying causal emergence in complex networks. Experiments were conducted on artificial networks as well as four types of real networks. In the case of ER random networks, the size of effective information was solely dependent on the connection probability $p$ and converged to $-\log_{2}p$ with the increase in network size. Additionally, a significant finding was the presence of a phase transition point, where the average degree of the network reached approximately $\log_{2}N$. Beyond this point, the random network structure did not contain additional information with increasing size. In the case of preferential attachment (PA) networks, when $\alpha<1.0$ ($\alpha$ represents the degree of preferential attachment), the effective information of the network increased as the network size expanded. Conversely, when $\alpha>1.0$, the opposite was observed. The scale-free network corresponding to $\alpha=1.0$ represented the critical boundary of growth. Regarding real networks, the authors found that biological networks exhibited the lowest EI due to the presence of significant noise, which can be removed through effective coarse-graining, and causal emergence was the most significant compared to other types of networks. However, technical networks exhibited sparsity and non-degeneracy, resulting in higher average efficiency and more specific node relationships. Consequently, they exhibited the highest EIs among the studied networks.
@@
The network coarse-graining method mentioned above is the greedy algorithm. However, when the network is very large, the efficiency of this method remains considerably low. Following that, Griebenow et al. Griebenow et al. (2019) introduced a spectral clustering-based approach for identifying causal emergence within networks. More specifically, the method involved performing eigenvalue decomposition of the TPM, followed by constructing a similarity matrix using the eigenvectors of the nodes. The OPTICS algorithm was employed to cluster the nodes, and nodes belonging to the same cluster were aggregated into a macro-node. Subsequently, the maximum value of EI was selected by utilizing the linear search distance hyperparameter $\epsilon$. In their paper, the authors additionally proposed a gradient descent algorithm based on deep learning. This approach encompassed several steps, beginning with the random initialization of a grouping matrix. Then, this matrix was used to construct a macro-network by combining the micro-networks. Finally, the grouping strategies were automatically learned through the process of maximizing effective information within the macro-network. However, this method often falls into the local optimal solution.